SOLVE WITH A FINANCIAL CALCULATOR
Note: You MUST show ALL your work to receive FULL credit-that is, show timelines and calculator sequence to justify your answers. He is entitled to receive one of the following prizes: Scott Frost has just won the Cornhusker Sweepstakes. a. $100,000 per year forever with the first payment received today b, $40,000 beginning next year growing by 4% forever. c. $200,000 for ten years starting in year 6. d. Ten payments of $150,000 per year starting one year from today. e. Five payments of $300,000 per year starting one year from today, then five more payments of $300,000 1. beginning eleven years from today. a discount rate of 10%, evaluate all of the alternative prizes and choose the prize that would have the greatest wealth maximizing effect for Frosty. (10 points)
Answer #1

1.
=100000/10%*1.1
=$1,100,000.00

2.
=40000/(0.1-0.04)
=$666,666.67

3.
N=10
PMT=200000
FV=0
I/Y=10%
CPT PV=$1,228,913.42

N=5
PMT=0
FV=1228913.42
I/Y=10%
CPT PV=$763,058.55

FInal PV=$763,058.55

4.
N=10
PMT=150000
FV=0
I/Y=10%
CPT PV=$921,685.07

5.
N=5
FV=0
PMT=300000
I/Y=10%
CPT PV=$1,137,236.03

N=10
PMT=0
FV=1137236.03
I/Y=10%
CPT PV=$438,453.72

Hence, Total PV=438453.72+1137236.03=$1,575,689.75


Hence, choose e)